# Theory-agnostic searches for non-gravitational modes in black hole ringdown

### v2

@@ Abstract
In any extension of General Relativity (GR), extra fundamental degrees of freedom couple to gravity. Besides deforming GR forecasts in a theory-dependent way, this coupling generically introduces extra modes in the gravitational-wave signal. We propose a novel theory-agnostic test of gravity to search for these nongravitational modes in black hole merger ringdown signals. To leading order in the GR deviations, their frequencies and damping times match those of a test scalar or vector field in a Kerr background, with only amplitudes and phases as free parameters. This test will be highly valuable for future detectors, which will achieve signal-to-noise ratios higher than 100 (and as high as 1000 for space-based detectors such as LISA). Such sensitivity will allow measurement of these modes with amplitude ratios as low as 0.05 for ground-based detectors (and as low as 0.008 for LISA), relative to the fundamental mode, enabling stringent agnostic constraints or detection of scalar/vector modes. By applying this test to GW150914, GW190521, and GW200129, we find that the current evidence for an extra mode is comparable to that for the first gravitational overtone, but its inclusion modifies the inferred remnant spin.
Introduction. The black hole (BH) spectroscopy program [1, 2, 3, 4] plays a prominent role in the landscape of strong-field tests of General Relativity (GR) [5, 6, 7, 8] and provides a unique method for examining the nature of compact remnants formed post-coalescence [9]. This program focuses on extracting the remnant quasinormal modes (QNMs) [10, 11, 12, 13] during the ringdown phase of a binary merger. In the context of linear perturbation theory, the signal $h(t)$ at intermediate times after the merger is represented by a superposition of the QNMs of the remnant [14], as it transitions towards a stationary configuration. Schematically,
@@
where $A_{i}$, $\phi_{i}$, $f_{i}$, $\tau_{i}$ are the amplitude, phase, frequency, and damping time of the $i$-th QNM. If the remnant is a BH, GR predicts that the infinite spectrum of QNMs is uniquely determined by its mass and spin $(M_{f},\chi_{f})$. This provides opportunities for conducting multiple null-hypothesis tests of gravity [15, 16] and investigating the nature of the remnant [17, 18, 19].
As suggested by Lovelock’s theorem [20, 21], an almost unavoidable ingredient of theories beyond GR is the presence of extra degrees of freedom nonminimally coupled to gravity [22, 6]. Examples are ubiquitous and include scalar fields in scalar-tensor theories and Horndeski’s gravity [23] (and their vector counterpart [24]), high-curvature corrections to GR that predict extra (pseudo)scalars and dilaton fields [25, 26, 27], Einstein-Aether [28] and Hořava–Lifshitz [29] gravity that postulate an extra timelike vector field, and massive gravity [30, 31] with both scalar and vector dynamical degrees of freedom (see [6] for a review of GR extensions and their field content). Effective extra degrees of freedom are also unavoidable in any approach that treats GR as the leading order term in an effective-field-theory expansion (e.g., [27]) and in low-energy effective string theories. These nonminimally coupled fields may modify the stationary BH solutions, leading to deviations from the Kerr metric, and/or modify the dynamics of the theory. In either case, two generic predictions are: i) a deformation of the Kerr QNMs,
@@
and ii) the existence of extra modes in the gravitational signal, that can be excited during the ringdown. This second option is due to the fact that the nonminimal coupling between new degrees of freedom and gravity results in coupled systems of linear perturbation equations, which act as a coupled set of oscillators [32, 33, 34]. It is therefore natural to split the ringdown signal (1) into two contributions,
@@
Standard tests of gravity based on the ringdown are rooted in the first line of Eq. (Theory-agnostic searches for non-gravitational modes in black hole ringdown). Namely, they are aimed at measuring $\delta f_{i}$ and $\delta\tau_{i}$ and check whether they are compatible with the null hypothesis [5]. This can be accomplished in two complementary ways, either through a theory-agnostic or theory-dependent method. However, both approaches have their own limitations. In the theory-agnostic approach one uses the parametrization (2), where $f_{i}^{\rm Kerr}$ and $\tau_{i}^{\rm Kerr}$ are known functions of the BH mass and spin, while $\delta f_{i}$ and $\delta\tau_{i}$ depend on the mass, spin, and all extra fundamental coupling constants of the theory. It is convenient (and natural, from an effective-field-theory perspective) to restrict to small GR deviations that continuously deform the Kerr result, in which case $\delta f_{i},\delta\tau_{i}\ll 1$ and are proportional to some combination of the mass and coupling constants. However, even in this case they are still generic functions of the spin $\chi_{f}$. Current parametrizations either neglect such spin dependence [5, 35, 36] or consider a small spin expansion of each deviation [37, 38, 39], which inevitably inflates the number of free parameters in the model. In the theory-dependent approach, the QNMs are computed in a given theory of gravity. For most theories this can be done again only in the small-coupling limit and very often perturbatively in the spin [40, 41, 32, 42, 43, 44, 45, 46, 47, 48, 49]. To reach good convergence, one needs to push the spin expansion to very high order [48], which is very challenging from the technical point of view. Alternatively, one needs to solve intricate systems of coupled partial differential equations [50, 51, 52, 53, 54, 55]. This approach has the benefit of limiting the number of free parameters to the sole coupling constants and BH spin, but must be performed on a case-by-case basis for every given theory.
@@
Given the above limitations, it would be highly desirable to develop complementary ringdown tests which are both theory-agnostic and accurate. In this work we explore a currently unbeaten path, related to the second line of Eq. (Theory-agnostic searches for non-gravitational modes in black hole ringdown). Namely, we propose to look for extra modes in the ringdown signal, which are not related to deformations of the Kerr ones. For concreteness, let us consider the case of an extra scalar degree of freedom nonminimally coupled to gravity (the same argument applies to other types of fields). Due to the coupling, the gravitational perturbations will contain also scalar modes (e.g., [41, 32, 56, 34] for two concrete examples in theories with quadratic curvature terms),
@@
where $f_{i}^{{\rm Kerr},\,s=0}$ and $\tau_{i}^{{\rm Kerr},\,s=0}$ are the QNMs of a test scalar field in the Kerr metric, and also in this case $\delta\hat{f}_{i}$ and $\delta\hat{\tau}_{i}$ are complicated, theory-dependent, functions of the mass, spin, and coupling constants. Crucially, in this case the amplitudes $\hat{A}_{i}$ of these modes are proportional to (powers of) the coupling constants [41, 32, 56, 34], and they must vanish in the GR limit. Therefore, to leading order in the corrections, we can neglect $\delta\hat{f}_{i}$ and $\delta\hat{\tau}_{i}$, so that the GR deviations are generically parametrized only by the amplitude of the test-field modes. This is precisely what happens in so-called dynamical Chern-Simons gravity [26, 32, 34] (see also Appendix), although it is a generic feature.
@@
where for simplicity we have neglected the GR deviations in the first line, since those are very well studied by standard ringdown tests (and subjected to the aforementioned limitations). In practice, here we will focus on a standard GR ringdown waveform (first line of Eq. (Theory-agnostic searches for non-gravitational modes in black hole ringdown)) augmented by new extra modes. Remarkably, to leading order these extra modes are known functions of the BH mass and spin, since they are those of a free test (scalar, vector, etc) field propagating on the Kerr metric (see, e.g., [57, 58] for tabulated values). This allows searching for extra modes in a theory-agnostic way, where the amplitudes and phases of the extra modes are the only beyond-GR parameters. In this sense, this test is reminiscent of searches for extra (scalar, vector) polarizations in GW signals in a theory-agnostic fashion [59, 60, 61, 5] and is complementary to ordinary ringdown tests (see, e.g., [62, 63, 5, 35, 36, 64, 65]), or to test with multiple free modes [5].
Searching for extra ringdown modes. The ringdown signal comprises of two polarizations and the modes are decomposed in a basis of spin-weighted spheroidal harmonics that depend on the remnant spin inclination angle [65]. Each mode is identified by three integers, $i\equiv(lmn)$, namely the multipolar, azimuthal, and overtone index, respectively, where $n=0$ corresponds to the fundamental tone and $(lmn)=(220)$ is the dominant mode in a non-precessing quasicircular coalescence. For concreteness, here we focus on the most interesting quadrupolar ($l=2$) case and neglect spin precession of the progenitor binary, but our test can be applied also to higher-order modes and to the precessing case (see Appendix). In particular, our ringdown waveform model has the following parameters:
@@
where $j={0,\dots,N}$, with $N$ the total number of overtones, whereas $\hat{A}_{220}^{s=0,1}$ and $\phi_{220}^{s=0,1}$ are the amplitude and phase of the extra scalar ($s=0$) or vector ($s=1$) $(220)$ mode. We will dub a model with $N$ tones and no extra mode GRN, whereas we will denote as GRN+S (GRN+V) a model with $N$ gravitational tones and an extra scalar (vector) mode. In practice, we define the amplitude ratio $A_{R}$ of a given mode relative to the gravitational fundamental one. We fix the luminosity distance $d_{L}$, the sky location of the system, and the inclination angle. Note that, for small-redshift sources (including those detected so far and presumably the loudest events detected in the future) $d_{L}$ is degenerate with the overall ringdown amplitude and can therefore be neglected without loss of generality. However, the possibility of neglecting inclination angle is a prerogative of our test, since it involves only $(l=m=2)$ modes, which have the same pattern functions and spheroidal-harmonic decomposition. Compared to common tests based on subleading modes of different $(l,m)$, this allows for the practical simplification of sampling in only one of these three degenerate parameters.
@@
The frequencies and damping times of the relevant modes are shown in the Appendix. At variance with overtones, the frequency of the $(220)$ scalar or vector mode is always well separated from that of the fundamental gravitational mode (and hence more easily resolvable from the latter), while the damping time is comparable (and hence the mode survives longer than overtones in the signal, almost as long as the fundamental gravitational mode).
@@
Bayesian analysis on real data. We exemplify our test on real events by performing a Bayesian parameter estimation using the PyCBC Inference code infrastructure [66]. The analysis aims to compute the posterior distribution of the parameters (6). We apply this test to three events: (i) GW150914 [67], the first GW event ever detected by LIGO (and still so far the one with the largest ringdown signal-to-noise ratio (SNR)), for which some debated evidence of overtones has been reported [15, 68, 69, 70, 71, 72, 73, 74, 75, 76, 35, 65]; (ii) GW190521 [77], a peculiar event in the upper mass gap for which a tentative detection of the $(330)$ and other modes (and possibly precession) has been obtained [78, 79] and which is prone also to ringdown amplitude-phase consistency tests [64]; (iii) GW200129, a peculiar loud event showing some tension with GR in some inspiral-merger-ringdown tests [5], tentatively ascribed to mismodelling of precession [80, 81]. As a proof of principle for the test, our parametrization assumes plane reflection symmetry, which is valid for spin-aligned progenitor binaries. In the Appendix we show that relaxing this assumption does not affect the results for GW190521 and GW200129. For the luminosity distance, inclination, and sky location we adopt the maximum likelihood values reported in [82].
@@
We use a gated-and-inpainted Gaussian likelihood noise model [83, 84, 78] to remove the influence of the pre-peak/non-ringdown times. The strain data within a time interval $t\in[t_{c}+t_{\rm offset}-0.5s,t_{c}+t_{\rm offset}]$ are replaced/inpainted such that the filtered inverse power spectral density is zero at all the times corresponding to the chosen interval [84]. Here, $t_{c}$ is the coalescence time, while $t_{\rm offset}$ defines the time in which we start our ringdown analysis.
@@
Figure 1: $\log_{10}$ Bayes factors for various ringdown models with extra scalar or vector modes with respect to the GR1 model as a function of the offset time $t_{\rm offset}$. Different colors and line styles denote different events, while different markers show the chosen model.
@@
Figure 2: Left panel: Minimum SNR necessary for detecting a scalar mode at $1-\sigma$ confidence level, according to the detectability criterion outlined in [85] using a Fisher matrix approximation. We take as a reference an event with remnant mass $M_{f}=62M_{\odot}$ for a range of amplitude ratios $A_{R,220}\in[0.02,1]$. The blue and red dashed curves denote a different remnant spin $\chi_{f}$ with the phase difference between the GR and scalar modes fixed to $\delta\phi=0$. The dependence on $\delta\phi\in[0,2\pi]$ is bracketed by the corresponding colored bands. The horizontal shaded bands represent the typical ringdown SNR expected for a GW150914-like event observed by the LVK network at design sensitivity (light blue), A+ (light green), ground-based third-generation observatories as ET/CE (yellow), and for a mass-rescaled version of GW150914 with $10^{5}M_{\odot}$ for LISA (orange). Right: The same quantity shown in a contour plot on the $(A_{R,220},\delta\phi)$ plane for fixed $\chi_{f}=0.67$.
@@
Our main results are summarized in Fig. 1, presenting the statistical evidence for different waveform models for these events and for different choices of $t_{\rm offset}$. We show $\log_{10}{\cal B}_{\texttt{GR1}}^{\rm Model}$, where the Bayes factor ${\cal B}_{\texttt{GR1}}^{\rm Model}$ is the ratio between the evidence of a given model and that of GR1 (i.e., a model containing only the fundamental gravitational mode and the first overtone). According to Jeffreys’ scale criterion [86], a $\log_{10}$ Bayes factor larger than $1$ (resp., $2$) would imply a strong (resp., decisive) Bayesian evidence in favor of a given model relative to GR1. The small values of $\log_{10}{\cal B}_{\texttt{GR1}}^{\rm Model}$ shown in Fig. 1 for any $t_{\rm offset}$ indicate that all models with extra scalar or vector modes have the same evidence as the GR1 one, presumably because the SNR in the ringdown for these events is not sufficiently high to exclude the presence of an extra scalar mode. This is consistent with what we shall discuss below with synthetic data.
@@
Table 1: 90% credible intervals for some of the parameters of event GW150914, assuming $t_{\rm offset}=2\,{\rm ms}$ (see Appendix for the posterior distributions). We denote the amplitude ratio of the scalar-to-tensor mode as $A_{R,220}=\hat{A}_{220}^{s=0}/A_{220}$.

### v3

@@ Abstract
In any extension of General Relativity (GR), extra fundamental degrees of freedom couple to gravity. Besides deforming GR forecasts in a theory-dependent way, this coupling generically introduces extra modes in the gravitational-wave signal. We propose a novel theory-agnostic test of gravity to search for these nongravitational modes in black hole merger ringdown signals. To leading order in the GR deviations, their frequencies and damping times match those of a test scalar or vector field in a Kerr background, with only amplitudes and phases as free parameters. By applying this test to GW150914, GW190521, and GW200129, we find no strong evidence for an extra mode; however, its inclusion modifies the inferred distribution of the remnant spin. This test will be applicable for future detectors, which will achieve signal-to-noise ratios higher than 100 (and as high as 1000 for space-based detectors such as LISA). Such sensitivity will allow measurement of these modes with amplitude ratios as low as 0.02 for ground-based detectors (and as low as 0.003 for LISA), relative to the fundamental mode, enabling stringent agnostic constraints or detection of scalar/vector modes.
I Introduction
The black hole (BH) spectroscopy program [Dreyer:2003bv, Detweiler:1980gk, Berti:2005ys, Gossan:2011ha] plays a prominent role in the landscape of strong-field tests of General Relativity (GR) [LIGOScientific:2021sio, Berti:2015itd, Berti:2018vdi, Colleoni:2024lpj] and provides a unique method for examining the nature of compact remnants formed post-coalescence [Cardoso:2019rvt]. This program focuses on extracting the remnant quasinormal modes (QNMs) [Vishveshwara:1970zz, Kokkotas:1999bd, Berti:2009kk, Konoplya:2011qq] during the ringdown phase of a binary merger. In the context of linear perturbation theory, the gravitational-wave (GW) signal $h(t)$ at intermediate times after the merger is represented by a superposition of the QNMs of the remnant [Leaver:1986gd], as it transitions towards a stationary configuration. Schematically,
@@ I Introduction
where $A_{i}$, $\phi_{i}$, $f_{i}$, $\tau_{i}$ are the amplitude, phase, frequency, and damping time of the $i$-th QNM, whereas $i\equiv(l,m,n)$ collectively represents the multipolar, azimuthal, and overtone index, respectively.
If the remnant is a BH, GR predicts that the infinite spectrum of QNMs is uniquely determined by its mass and spin $(M_{f},\chi_{f})$. This provides opportunities for conducting multiple null-hypothesis tests of gravity [Isi:2019aib, Franchini:2023eda] and investigating the nature of the remnant [Maggio:2020jml, Maggio:2021ans, Maggio:2023fwy].
As suggested by Lovelock’s theorem [Lovelock:1971yv, Lovelock:1972vz], an almost unavoidable ingredient of theories beyond GR is the presence of extra degrees of freedom nonminimally coupled to gravity [Sotiriou:2014yhm, Berti:2015itd]. Examples are ubiquitous and include scalar fields in scalar-tensor theories and Horndeski’s gravity [Horndeski:1974wa] (and their vector counterpart [Heisenberg:2014rta]), high-curvature corrections to GR that predict extra (pseudo)scalars and dilaton fields [Kanti:1995vq, Alexander:2009tp, Endlich:2017tqa], Einstein-Aether [Jacobson:2000xp] and Hořava–Lifshitz [Horava:2009uw] gravity that postulate an extra timelike vector field, and massive gravity [deRham:2010kj, Hassan:2011zd] with both scalar and vector dynamical degrees of freedom (see [Berti:2015itd] for a review of GR extensions and their field content). Effective extra degrees of freedom are also unavoidable in any approach that treats GR as the leading order term in an effective-field-theory expansion (e.g., [Endlich:2017tqa]) and in low-energy effective string theories. These nonminimally coupled fields may modify the stationary BH solutions, leading to deviations from the Kerr metric, and/or modify the dynamics of the theory. In either case, two generic predictions are: i) a deformation of the Kerr QNMs,
@@
and ii) the existence of extra modes in the gravitational signal, that can be excited during the ringdown. This second option is due to the fact that the nonminimal coupling between new degrees of freedom and gravity results in coupled systems of linear perturbation equations, which act as a coupled set of oscillators [Molina:2010fb, Pani:2013pma, Cardoso:2020nst, DAddario:2023erc]. It is therefore natural to split the ringdown signal (1) into two contributions,
@@
Standard tests of gravity based on the ringdown are rooted in the first line of Eq. (I). Namely, they are aimed at measuring $\delta f_{i}$ and $\delta\tau_{i}$ and check whether they are compatible with the null hypothesis [LIGOScientific:2021sio]. This can be accomplished in two complementary ways, either through a theory-agnostic or theory-dependent method. However, both approaches have their own limitations. In the theory-agnostic approach one uses the parametrization (2), where $f_{i}^{\rm Kerr}$ and $\tau_{i}^{\rm Kerr}$ are known functions of the BH mass and spin, while $\delta f_{i}$ and $\delta\tau_{i}$ depend on the mass, spin, and all extra fundamental coupling constants of the theory. We will focus on small deviations from GR. This is motivated by an effective-field-theory perspective wherein Einstein’s theory is the lowest term in an expansion containing all possible higher-order operators, and also by the stringent observational constraints already placed by GW observations. As a result, any deviation from the standard GR prediction must be small and proportional to (powers of) the coupling constants of the theory. In particular, $\delta f_{i},\delta\tau_{i}\ll 1$ and are proportional to some combination of the mass and coupling constants. However, even in this case they are still generic, theory-dependent functions of the spin $\chi_{f}$. Current parametrizations either neglect such spin dependence [LIGOScientific:2021sio, Ma:2023cwe, Ma:2022wpv] or consider a small spin expansion of each deviation [Maselli:2019mjd, Carullo:2021dui, Maselli:2023khq], which inevitably inflates the number of free parameters in the model. In the theory-dependent approach, the QNMs are computed in a given theory of gravity. For most theories this can be done again only in the small-coupling limit and very often perturbatively in the spin [Pani:2009wy, Cardoso:2009pk, Molina:2010fb, Pani:2013ija, Pani:2013wsa, Pierini:2021jxd, Wagle:2021tam, Cano:2021myl, Pierini:2022eim, Cano:2023jbk, Wagle:2023fwl]. To reach good convergence, one needs to push the spin expansion to very high order [Cano:2023jbk], which is very challenging from the technical point of view. Alternatively, one needs to solve intricate systems of coupled partial differential equations [Dias:2015wqa, Chung:2023zdq, Chung:2023wkd, Blazquez-Salcedo:2023hwg, Chung:2024ira, Chung:2024vaf]. This approach has the benefit of limiting the number of free parameters to the sole coupling constants and BH spin, but must be performed on a case-by-case basis for every given theory.
Given the above limitations, it would be highly desirable to develop complementary ringdown tests which are both theory-agnostic and accurate. In this work we explore a currently unbeaten path, related to the second line of Eq. (I). Namely, we propose to look for extra modes in the ringdown signal. Note that these extra modes are unavoidably present in beyond-GR theories, raising the important issue that current ringdown analyses (based only on the first line of Eq. (I)) are incomplete.
@@
For concreteness, let us consider the case of an extra scalar degree of freedom nonminimally coupled to gravity (the same argument applies to other types of fields). Due to the coupling, the gravitational perturbations will contain also scalar modes (e.g., [Cardoso:2009pk, Molina:2010fb, Blazquez-Salcedo:2016enn, Cardoso:2020nst] for two concrete examples in theories with quadratic curvature terms),
@@
where $f_{i}^{{\rm Kerr},\,s=0}$ and $\tau_{i}^{{\rm Kerr},\,s=0}$ are the QNMs of a test scalar field in the Kerr metric, and also in this case $\delta\hat{f}_{i}$ and $\delta\hat{\tau}_{i}$ are complicated, theory-dependent, functions of the mass, spin, and coupling constants, which incorporate both deviations from the GR BH background and modified dynamics. Crucially, in this case the amplitudes $\hat{A}_{i}$ of these modes are proportional to (powers of) the coupling constants [Cardoso:2009pk, Molina:2010fb, Blazquez-Salcedo:2016enn, Cardoso:2020nst, DAddario:2023erc], and they must vanish in the GR limit. Therefore, to leading order in the corrections, we can neglect $\delta\hat{f}_{i}$ and $\delta\hat{\tau}_{i}$, so that the GR deviations are generically parametrized only by the amplitude of the test-field modes on a GR BH background. This is precisely what happens in so-called dynamical Chern-Simons gravity [Alexander:2009tp, Molina:2010fb, Cardoso:2020nst] (see also Appendix VII.1), although it is a generic feature [DAddario:2023erc].
@@
where for simplicity we have neglected the GR deviations in the first line, since those are very well studied by standard ringdown tests (and subjected to the aforementioned limitations). Notably, as explained below and although our framework is generic, one of its advantages is that it can be applied to a single multipole couple $(l,m)$, in which case $i=0,1,2,..$ would simply correspond to the overtone number $n=0,1,2,..$ at fixed $(l,m)$. If one includes only the dominant GR fundamental mode in the analysis ($i\equiv(2,2,0)$ in the first line of Eq. (I)), as we will do below for the so-called GR0+S or GR0+V models, the corrections $\delta f_{220}$ and $\delta\tau_{220}$ to the fundamental gravitational QNM are degenerate with the final mass and spin. Therefore, they can be neglected without loss of generality. Instead, including also gravitational overtones would require adding extra corrections $\delta f_{i}$ and $\delta\tau_{i}$ (with $i\equiv(2,2,n\geq 1)$) in the first line of Eq. (I), as typically done in ordinary BH spectroscopy tests with overtones. In the models considered here, the mentioned degeneracy is broken as we neglect higher harmonics but include at least two modes through either an overtone (GR1 model) or an extra scalar/vector mode ($\texttt{GR0+S}/\texttt{GR0+V}$) in addition to the fundamental GR mode, determining the final mass and spin uniquely. Explicitly, our models contain the modes: $\texttt{GR1}\rightarrow(2,2,0),(2,2,1)$; $\texttt{GR0+S}\rightarrow(2,2,0),(2,2,0)_{\text{scalar}}$; $\texttt{GR0+V}\rightarrow(2,2,0),(2,2,0)_{\text{vector}}$.
In practice, here we will focus on a standard GR ringdown waveform (first line of Eq. (I)) augmented by new extra modes. Remarkably, to leading order these extra modes are known functions of the BH mass and spin, since they are those of a free test (scalar, vector, etc) field propagating on the Kerr metric (see, e.g., [BertiWebpage, CardosoWebpage] for tabulated values). This allows searching for extra modes in a theory-agnostic way, where the amplitudes and phases of the extra modes are the only beyond-GR parameters. In this sense, this test is reminiscent of searches for extra (scalar, vector) polarizations in GW signals in a theory-agnostic fashion [Takeda:2018uai, Takeda:2019gwk, Isi:2017fbj, LIGOScientific:2021sio] and is complementary to ordinary ringdown tests (see, e.g., [Brito:2018rfr, Carullo:2019flw, LIGOScientific:2021sio, Ma:2023cwe, Ma:2022wpv, Forteza:2022tgq, Baibhav:2023clw]), or to test with multiple free modes [LIGOScientific:2021sio].
II Searching for extra ringdown modes
The ringdown signal comprises of two polarizations and the modes are decomposed in a basis of spin-weighted spheroidal harmonics that depend on the remnant spin inclination angle [Baibhav:2023clw]. In a non-precession quasicircular coalescence, $(lmn)=(220)$ is the dominant mode. For concreteness, here we focus on the most interesting quadrupolar ($l=2$) case and neglect spin precession of the progenitor binary, but our test can be applied also to higher-order modes and to the precessing case (see Appendix VII.2). In particular, our ringdown waveform model has the following parameters:
@@ II Searching for extra ringdown modes
where $j={0,\dots,N}$, with $N$ the total number of overtones, whereas $\hat{A}_{220}^{s=0,1}$ and $\phi_{220}^{s=0,1}$ are the amplitude and phase of the extra scalar ($s=0$) or vector ($s=1$) $(220)$ mode. We will dub a model with $N$ tones and no extra mode GRN, whereas we will denote as GRN+S (GRN+V) a model with $N$ gravitational tones and an extra scalar (vector) mode. In practice, we define the amplitude ratio $A_{R}$ of a given mode relative to the gravitational fundamental one. Notice that in general, these amplitudes depend on the specific theory (see an example for dynamical Chern-Simons theory in the Appendix VII.1), and estimating their value in a merger requires performing nonlinear coalescence simulations beyond GR. For this reason, it is not easy to predict the relevant importance of extra modes with respect to other GR effects, e.g., overtones or nonlinearities. For the sake of generality, here we remain agnostic on the amplitude modeling. We fix the luminosity distance $d_{L}$, the sky location of the system, and the inclination angle as done standard ringdown analysis [LIGOScientific:2021sio, Carullo:2021dui, Isi:2019aib, Capano:2021etf]. Although marginalization would be the ideal approach, it involves its own challenges, like the accurate excision of the pre-merger part to avoid the contamination of the post-merger data. Failing on the accurate excision could impact negatively on the accuracy of the no-hair theorem tests performed using the post-merger data [Correia:2023bfn, Correia:2023ipz]. Furthermore, the sky-marginalization has been shown to introduce only mild differences for current GW events, from [Finch:2022ynt, Correia:2023ipz] where the results obtained, are still consistent with those of [Carullo:2021dui, Isi:2019aib] which fixed the sky location to the maximum likelihood values. However, the scenario might be different for higher-SNR events and with the inclusion of higher harmonics. Note that, for small-redshift sources (including those detected so far and presumably the loudest events detected in the future) $d_{L}$ is degenerate with the overall ringdown amplitude and can therefore be neglected without loss of generality. However, the possibility of neglecting inclination angle is a prerogative of our test, since it can involve only $(l=m=2)$ modes , which have the same pattern functions and spheroidal-harmonic decomposition (similarly to overtone-based tests but without the issues related to overtones). Compared to common tests based on subleading modes of different $(l,m)$, this allows for the practical simplification of sampling in only one of these three degenerate parameters. The test can be expanded to include higher harmonics, in which case one would have also to include the inclination as an extra parameter.
@@
The frequencies and damping times of the relevant modes are shown in the Appendix VII.3. At variance with overtones, the frequency of the $(220)$ scalar or vector mode is always well separated [^1] Except possibly in the $\chi_{f}\rightarrow 1$ limit, which is not relevant for the spin values considered here. from that of the fundamental gravitational mode (and hence more easily resolvable from the latter), while the damping time is comparable (and hence the mode survives longer than overtones in the signal, almost as long as the fundamental gravitational mode).
III Bayesian analysis on real data
@@
We exemplify our test on real events by performing a Bayesian parameter estimation using the PyCBC Inference code infrastructure [Biwer:2018osg]. The analysis aims to compute the posterior distribution of the parameters (6). We apply this test to three events: (i) GW150914 [LIGOScientific:2016aoc], the first GW event ever detected by LIGO (and still so far the one with the largest ringdown signal-to-noise ratio (SNR)), for which some debated evidence of overtones has been reported [Isi:2019aib, Giesler:2019uxc, Ota:2019bzl, Bhagwat:2019dtm, CalderonBustillo:2020rmh, Cotesta:2022pci, Isi:2022mhy, Finch:2022ynt, Carullo:2023gtf, Ma:2023vvr, Ma:2023cwe, Baibhav:2023clw]; (ii) GW190521 [LIGOScientific:2020iuh], a peculiar event in the upper mass gap for which a tentative detection of the $(330)$ and other modes (and possibly precession) has been obtained [Capano:2021etf, Siegel:2023lxl] and which is prone also to ringdown amplitude-phase consistency tests [Forteza:2022tgq]; (iii) GW200129, a peculiar loud event showing some tension with GR in some inspiral-merger-ringdown tests [LIGOScientific:2021sio], tentatively ascribed to mismodelling of precession [Maggio:2022hre, Gupta:2024gun]. As a proof of principle for the test, our parametrization assumes plane reflection symmetry, which is valid for spin-aligned progenitor binaries. In the Appendix VII.2, we show that relaxing this assumption does not affect the results for GW190521 and GW200129. For the luminosity distance, inclination, and sky location we adopt the maximum likelihood values reported in [Nitz:2021zwj].
@@
We use a gated-and-inpainted Gaussian likelihood noise model [Usman:2015kfa, Zackay:2019kkv, Capano:2021etf] to remove the influence of the pre-peak/non-ringdown times. The strain data within a time interval $t\in[t_{c}+t_{\rm offset}-0.5s,t_{c}+t_{\rm offset}]$ are replaced/inpainted such that the filtered inverse power spectral density is zero at all the times corresponding to the chosen interval [Zackay:2019kkv]. Here, $t_{c}$ is the coalescence time, while $t_{\rm offset}$ defines the time in which we start our ringdown analysis.
@@
Figure 1: $\log_{10}$ Bayes factors for various ringdown models with extra scalar or vector modes (labelled with ’A’) with respect to the GR1 model as a function of the offset time $t_{\rm offset}$. Different colors and line styles denote different events, while different markers show the chosen model. Note that each event has a different mass thus, a different time scale $t_{\text{offset}}$. Therefore, the $\log_{10}$ Bayes factor trends, specially at negative times, are expected to differ.
@@
Figure 2: Left panel: Minimum SNR necessary for detecting a scalar mode at $1-\sigma$ confidence level, according to the detectability criterion outlined in [Berti:2005ys, JimenezForteza:2020cve] using a Fisher matrix approximation, which makes our estimates approximately independent of the sensitivity-curve profile. The blue and red dashed curves denote a different remnant spin $\chi_{f}$ with the phase difference between the GR and scalar modes fixed to $\delta\phi=0$. The dependence on $\delta\phi\in[0,2\pi]$ is bracketed by the corresponding colored bands. The horizontal shaded bands represent the expected ringdown SNR for a GW150914-like event with a remnant mass of $M_{f}=62\,M_{\odot}$ as observed by ground-based detectors, including the LIGO-Virgo-KAGRA network (light blue), A+ (light green), and CE/ET (yellow). The orange band corresponds to the expected ringdown SNR for a similar event with a remnant mass of $M_{f}=10^{7}\,M_{\odot}$ observed by LISA (orange). We consider the range of amplitude ratios $A_{R,220}\in[0.001,1]$. Right: The same quantity shown in a contour plot on the $(A_{R,220},\delta\phi)$ plane for fixed $\chi_{f}=0.67$.
@@
Our main results are summarized in Fig. 1, presenting the statistical evidence for different waveform models for these events and for different choices of $t_{\rm offset}$. We show $\log_{10}{\cal B}_{\texttt{GR1}}^{\rm A}$, where the Bayes factor ${\cal B}_{\texttt{GR1}}^{\rm A}$ is the ratio between the evidence of a given model (’A’) and that of GR1 (i.e., a model containing only the fundamental gravitational mode and the first overtone). According to Jeffreys’ scale criterion [Jeffreys:1939xee, DelPozzo:2011pg, Gossan:2011ha, CalderonBustillo:2020rmh], a $\log_{10}$ Bayes factor larger than $1$ (resp., $2$) would imply a strong (resp., decisive) Bayesian evidence in favor of a given model relative to GR1. The small values of $\log_{10}{\cal B}_{\texttt{GR1}}^{\rm A}$ shown in Fig. 1 for any $t_{\rm offset}$ indicate that all models with extra scalar or vector modes have the same evidence as the GR1 one, presumably because the SNR in the ringdown for these events is not sufficiently high to exclude the presence of an extra scalar mode. This is consistent with what we shall discuss below with synthetic data.
@@
Table 1: 90% credible intervals for some of the parameters of event GW150914, assuming $t_{\rm offset}=2\,{\rm ms}$ (see Appendix VII.3 for the posterior distributions).